The market demand and supply for renting apartments in the Rosebank areas are given as: Demand:P=12 000-5Q. Supply: P=800+3Q

Calculate the equilibrium quantity and price


1
Expert's answer
2021-06-06T21:45:03-0400

"Q_D=Q_S"


"12,000-5Q=800+3Q"

"12,000-800=5Q+3Q"

"11,200=8Q"

"Q=1,400"


"P=12,000-5Q"

"P=12,000-5(1400)"

"P=12,000-7000"

"P=5,000"